Two dice were rolled. Answer the questions:

I) What is the probability that the first die will have a number that is a multiple of 2 as the outcome? (10 points)

a. 1/2

b. 1/5

c. 1/6

d. 1/36

e. None of the above

II) What is the probability that the first die will have “a number less than 4” as the outcome? (10 points)

a. 1/5

b. 1/3

c. 1/6

d. 1/36

e. None of the above

III) What is the probability that the sum of the two dice is divisible by 11? (10 points)

a. 1/18

b. 1/12

c. 1/36

d. 1/9

e. None of the above
Answer #1
(1,1)
(1,2)(2,1)
(1,3)(3,1)(2,2)
(1,4)(4,1)(2,3)(3,2)
(1,5)(5,1)(2,4)(4,2)(3,3)
(1,6)(6,1)(2,5)(5,2)(3,4)(4,3)
(2,6)(6,2)(3,5)(5,3),(4,4)
(3,6)(6,3)(4,5)(5,4)
(4,6)(6,4)(5,5)
(5,6)(6,5)
(6,6)

1)

Probability = 18/36 = 1/2

a. 1/2

2)

Probability = 18/36 = 1/2

e. None of the above

3)

(6,5) (5,6) possible out of 36 cases

Probability = 2/ 36

= 1/18

a. 1/18


THANKS

revert back for doubt

please upvote